10x^4-9x^2+2=0
I set x^2=u
I got u=2/5
u=1/2
when I use the square root property x^2= ±√(2/5) and x^2=±√(1/2), and the math lab says is wrong. Why???

Answer by richwmiller(17219)

x^2=u
10u^2-9u+2=0
x^2=2/5 not +-sqrt(2/5)
[x=+-sqrt(2/5) we don't need to know this; we never use it.]
x^2=1/2 not +-sqrt(1/2)
[x=+-sqrt(1/2) we don't need to know this; we never use it.]
check
10*1/4-9* 1/2+2=0
2.5-4.5+2=0
0=0
ok
check
10*4/25-9*2/5+2=0
40/25-18/5+2=0
8/5-18/5+10/5=0
0=0
ok